# Linearly Polarized Light-Induced Anomalous Hall Effect and Topological Phase Transitions in an Altermagnetic Topological Insulator

### 06 Mar 2026 v1

@@ III.3 LPL-induced topological phase transitions
In this section, we evaluate the topological phase transitions of out-of-plane $d$-wave AM and AFM under LPL irradiation. The topological phase transitions are usually accompanied by the closing and reopening of the band gap. We first calculate the evolution of the band gaps for spin-up and spin-down bands as a function of LPL intensity for AM and AFM that are initially in trivial or QSH phases, respectively. As shown in Figs. 4(b) and (d), the gaps for spin-up and spin-down bands remain identical since the AFM retains $\mathcal{PT}$ symmetry under LPL irradiation. In contrast, there is no symmetry linking the spin-up and spin-down channels once the LPL is applied to the AM. As a result, the evolution of the spin-up band gap differs from that of the spin-down band gap as the light intensity changes, as illustrated in Figs. 4(a) and (c).
If the material is topologically trivial in the absence of light, the energy gaps for both the AM and AFM configurations do not close with increasing light intensity, which are shown in Figs. 4(a) and (b), thus preventing a topological phase transition. For the AFM QSH insulator, as shown in Fig. 4(d), the band gaps of both spin channels close simultaneously with increasing light intensity, leading to concurrent transitions of $|C_{\uparrow}|$ and $|C_{\downarrow}|$ from 1 to 0. As a result, the QSH phase becomes a topologically trivial phase. Intriguingly, the AM QSH insulator exhibits a different evolution. As illustrated in Fig. 4(c), the spin-down band gap closes first upon increasing light intensity, driving $C_{\downarrow}$ from 1 to 0 while $C_{\uparrow}$ remains at 1-1. Consequently, the total Chern number becomes 1-1, indicating the emergence of a Chern insulator phase. With further increase of the LPL intensity, the remaining spin-up band gap also closes, eventually driving the material into a trivial phase. Therefore, the AM QSH insulator undergoes a sequential topological transition from the QSH phase to a Chern insulator phase, and finally to a trivial phase. Notably, the nonzero Chern number in this intermediate regime originates entirely from the spin-up bands, identifying it as a spin-polarized Chern insulator.
@@ IV Conclusion
We demonstrate that LPL breaks the $C_{4}\mathcal{T}$ symmetry connecting the spin-up and spin-down states in AM, while still preserving the $\mathcal{P}\mathcal{T}$ symmetry in AFM. First, within the high-frequency approximation, LPL lifts the spin degeneracy along high-symmetry lines in the effective Hamiltonian of the AM, driving it into a compensated ferrimagnetic state, whereas spin degeneracy remains preserved throughout the entire BZ in the AFM case. Secondly, LPL can induce an AHE in AM but fails to do so in AFM. Thirdly, LPL can drive a AM QSH insulator through a topological phase transition into a spin-polarized Chern insulator phase, thereby generating a spin-polarized QAHE. In contrast, LPL cannot induce a nonzero Chern number in a QSH AFM. Furthermore, we observe that the AHE becomes strongly anisotropic with respect to the polarization direction of the LPL, enabling continuous tuning and sign reversal upon polarization rotation. Our findings provide a perspective on the Floquet engineering of AM, as well as a potential method for experimentally distinguishing between AFM and AM materials and probing the spin-splitting pattern of AM.

### 13 Mar 2026 v2

@@ III.3 LPL-induced topological phase transitions
In this section, we evaluate the topological phase transitions of the $d$-wave AM and AFM under LPL irradiation. The topological phase transitions are usually accompanied by the closing and reopening of the band gap. We first calculate the evolution of the band gaps for spin-up and spin-down bands as a function of LPL intensity for AM and AFM that are initially in trivial or QSH phases, respectively. As shown in Figs. 4(b) and (d), the gaps for spin-up and spin-down bands remain identical since the AFM retains $\mathcal{PT}$ symmetry under LPL irradiation. In contrast, there is no symmetry linking the spin-up and spin-down channels once the LPL is applied to the AM. As a result, the evolution of the spin-up band gap differs from that of the spin-down band gap as the light intensity changes, as illustrated in Figs. 4(a) and (c).
If the material is topologically trivial in the absence of light, the energy gaps for both the AM and AFM configurations do not close with increasing light intensity, as shown in Figs. 4(a) and (b), thus preventing a topological phase transition. For the AFM QSH insulator, as shown in Fig. 4(d), the band gaps of both spin channels close simultaneously with increasing light intensity, leading to concurrent transitions of $|C_{\uparrow}|$ and $|C_{\downarrow}|$ from 1 to 0. As a result, the QSH phase becomes a topologically trivial phase. Intriguingly, the AM QSH insulator exhibits a different evolution. As illustrated in Fig. 4(c), the spin-down band gap closes first upon increasing light intensity, driving $C_{\downarrow}$ from 1 to 0 while $C_{\uparrow}$ remains at 1-1. Consequently, the total Chern number becomes 1-1, indicating the emergence of a Chern insulator phase. With further increase of the LPL intensity, the remaining spin-up band gap also closes, eventually driving the material into a trivial phase. Therefore, the AM QSH insulator undergoes a sequential topological transition from the QSH phase to a Chern insulator phase, and finally to a trivial phase. Notably, the nonzero Chern number in this intermediate regime originates entirely from the spin-up bands, identifying it as a spin-polarized Chern insulator.
@@ IV Conclusion
We demonstrate that LPL breaks the $C_{4z}\mathcal{T}$ symmetry connecting the spin-up and spin-down states in AM, while still preserving the $\mathcal{P}\mathcal{T}$ symmetry in AFM. First, within the high-frequency approximation, LPL lifts the spin degeneracy along high-symmetry lines in the effective Hamiltonian of the AM, driving it into a compensated ferrimagnetic state, whereas spin degeneracy remains preserved throughout the entire BZ in the AFM case. Secondly, LPL can induce an AHE in AM but fails in AFM. Thirdly, LPL can drive an AM QSH insulator into a spin-polarized Chern insulator through a topological phase transition, thereby generating a spin-polarized QAH effect. In contrast, LPL cannot induce a nonzero Chern number in a QSH AFM. Furthermore, we observe that the AHE becomes strongly anisotropic with respect to the polarization direction of the LPL, enabling continuous tuning and sign reversal upon polarization rotation. Our findings provide a perspective on the Floquet engineering of AM, as well as a potential method for experimentally distinguishing between AFM and AM materials and probing the spin-splitting pattern of AM.